# Deep generative modelling of canonical ensemble with differentiable thermal properties

### v1

@@ Abstract
Furthermore, we compared the mean energy and heat capacity obtained by the MCMC simulation and the differentiation of the free energy in Fig. 2(d). In comparison to the MCMC result, we observed only minor deviations in the heat capacity around the PT, indicating a good fit of the free energy to the second-order derivative and a good capture of the subtle PT transition. Numerical results in appendix also show that the derivative estimation gives better convergence than the statistical averaging [17]. In contrast to the one-by-one generation process of PixelCNN, the NF model is significantly faster due to its all-to-all style of generation. This leads to a comparable training-plus-evaluation time (36 hours on one Nvidia RTX 4090) to performing an MCMC sweep over many temperature points. Additionally, the low STD allows us to estimate the heat capacity using the first-order derivative, which offers a significant speed-up and reduces the memory requirement. We give a detailed description of this scheme in appendix [17].
Outlooks.
@@
In this Letter, we present a novel framework of variational training for the explicit density generative models. By converting the Boltzmann distribution optimization problem into a temperature-dependent loss function, we achieve a direct-sampling Boltzmann distribution model with explicit temperature dependence. This allows thermodynamic quantities, such as free energy and partition function, to be differentiable functions of temperature. Our proposed method is as accurate as the MCMC method, but more efficient due to direct sampling. Moreover, as a single model with temperature dependence, its direct-sampling estimation remains valid across a continuous temperature range. In contrast to previous deep generative models that optimize at a single temperature, our method preserves the essential temperature dependence akin to an analytical solution. As a result, there is no need for training additional models at discrete temperature points, and a single model can reliably capture subtle thermal effects, even at PTs.
@@
With minor modifications, we expect to have immediate practical applications by adapting our method to these popular deep generative variational models of single temperature [12, 13, 14, 15, 39, 40, 16]. Our proposed method can be seen as a generalization of the variational mean-field method, which consists of multiple conditional parameters. While our current method focuses on temperature, it opens up intriguing possibilities for exploring other parameters, such as the pressure and external magnetization fields. Here we mainly applied our method to lattice models, but it is also viable to extend its application to more realistic systems with many meta-stable states, e.g., atomisitic simulations.
@@ Acknowledgements.
We thank Lei Wang, Wen-Qing Xie, Chu Li, Tao Li, and Jun-Ting Yu for many useful discussions. S-H.L. acknowledges support from Hong Kong Research Grants Council (GRF-16302423). D.P. acknowledges support from the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (GRF-16301723), National Natural Science Foundation of China through the Excellent Young Scientists Fund (22022310), and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083). Part of this work was carried out using computational resources from the National Supercomputer Center in Guangzhou, China, and the X-GPU cluster supported by the HKRGC Collaborative Research Fund C6021-19EF.
Supplemental Information: Deep generative modelling of canonical ensemble with differentiable thermal properties
Appendix A Variational Minimization of the Free Energy
Given a probability distribution $q(\mathbf{x})$ and an energy function $E(\mathbf{x})$, the free energy is
$F=TS+\langle E\rangle=\sum_{\{\mathbf{x}\}}\frac{1}{\beta}q(\mathbf{x})\log q(\mathbf{x})+q(\mathbf{x})E(\mathbf{x}).$ (S1)
@@ Appendix A Variational Minimization of the Free Energy
Considering a $q^{*}$ that minimizes the free energy for all temperatures, then this $q^{*}$ should be the solution to the following optimization problem.
@@
This optimization can be analytically solved via the Lagrange multiplier method [41], which gives the Boltzmann distribution as a function of temperature, i.e.,
@@
Using an explicit density model, one can directly sample $\bm{x}$ to perform statistical averaging as an estimation of the summation over $\{\bm{x}\}$, and the probability normalization condition is automatically satisfied, so we get
$\min_{\theta}\mathop{\mathbb{E}}_{\beta\sim U}\frac{1}{\beta}\Big{[}\mathop{\mathbb{E}}_{\bm{x}\sim q_{\theta}(\cdot,\beta)}q_{\theta}(\mathbf{x})\log q_{\theta}(\mathbf{x})+\beta q_{\theta}(\mathbf{x})E(\mathbf{x})\Big{]},$ (S7)
@@
Appendix B Reparameterization and Reinforce Estimation
B.1 First-order and second-order reparameterization
The reparameterization introduces a new random variable $\epsilon$ and rewrites the original expectation into an expectation over the distribution of $\epsilon$, i.e.,
@@ B.1 First-order and second-order reparameterization
The derivatives can be reformulated as
@@
B.2 First-order and second-order reinforce estimation
@@
The variance of estimate of Eq. (S11) can be reduced via [23]
@@
where $b$ is the baseline. $b$ has many different forms [23] and here we use the simplest one:
@@
+$+\sum_{\{\mathbf{x}\}}q_{\theta}(\mathbf{x})\nabla_{\theta}^{2}f(\mathbf{x};\theta)+2\nabla_{\theta}f(\mathbf{x};\theta)q_{\theta}(\mathbf{x})\nabla_{\theta}\log q_{\theta}(\mathbf{x})$
@@
The variance reduction method can be also employed
@@
B.3 Reparameterization of acceptance-rejection sampling
The acceptance-rejection sampling is a common way to draw samples from certain relatively complex distributions. The reparameterization of acceptance-rejection sampling is given in Ref. [38]. For a better illustration, we first give the general framework of acceptance-rejection sampling.
In acceptance-rejection sampling, to sample distribution $q_{\theta}(z)$, we first draw a random variable $\epsilon$ from the distribution $s(\epsilon)$. It is then transformed using a parameterized transformation $h_{\theta}$, similar to the NF case, we should consider the Jacobian of the transformation. Then its possibility changes to $r_{\theta}({z})$, where ${z}=h_{\theta}(\epsilon)$. One accepts the sample with the probability $\min\{1,\frac{q_{\theta}({z})}{M_{\theta}r_{\theta}({z})}\}$ or rejects it, as in Alg. 1.
@@ 5       $u\sim{U}(0,1)$;
[…]
@@
The first-order derivative can be reformulated as
@@
Then, using the log-derivative trick, this becomes
@@
We generalize the scheme in [38], and give the second-order derivative
@@
This is
[…]
@@
[…]
so the variance reduction method [23] can also be applied here.
@@
Similar to the NF model case, the $r_{\theta}$ distribution can be derived using the $s$ distribution and the Jacobian of $h_{\theta}$.
Appendix C Macroscopic Behaviors from the Differentiation of the Free Energy
C.1 Analysis of the differentiation method compared to the statistical averaging with direct sampling
After training a direct-sample model with differentiable temperature condition, to estimate mean energy and heat capacity, other than the differentiation method in Eq. (2), one can also perform statistical averaging with direct sampling, i.e.,
@@ C.1 Analysis of the differentiation method compared to the statistical averaging with direct sampling
For a better analysis, We establish a connection between the two methods.
Using Eq. (2), given the estimated $-\log\bar{\mathcal{Z}}(\beta)$, the mean energy is
@@
A similar analysis can be done for the heat capacity with the same conclusion; that is, when the training is perfectly done, the differentiation method gives the same results as direct sampling. However, for imperfect training, the two methods give different results. The differentiation results, as they’re closely related to the minimization goal, have a better accuracy than the statistical average. Additionally, one can view the VaTD training as a generalization of the variational mean-field method. In the mean-field theory, when the variational ansatz $q_{\theta}$ is not flexible enough, the $q_{\theta}$ may not converge to the target distribution $p$, but still could have a relative accurate estimation of the partition function (or free energy).
@@
On the other hand, even with perfect or near-perfect training, it is better to choose the differentiation method, because the differentiation method has a much faster convergence speed. From Eq. (4), one can see that a perfect model will give uniformly the same estimate of the free energy for any $\bm{x}$. In this sense, a batch of just one sample will converge.
@@
The numerical tests in the following sections demonstrate these two advantages of the differentiation method.
C.2 Estimation of the heat capacity using first-order differentiation
As the second-order derivatives are not memory-efficient for backpropagation automatic differentiation, we propose a scheme to estimate the heat capacity using the first-order differentiation when the loss of the model is low enough.
Consider the first-order differentiation

### v2

@@ Abstract
Furthermore, in Fig. 4, we compared the mean energy and heat capacity obtained by the MCMC simulation and the differentiation of the estimated free energy. In comparison to the MCMC result, we observed only minor deviations around the PT, indicating a good fit of the free energy to the second-order derivative and a good capture of the subtle PT transition. Numerical results in SI also show that this free energy derivative estimation of mean energy and heat capacity gives better convergence than the statistical averaging with direct sampling [SM]. Moreover, the derivative approach expresses heat capacity as a mean value over a batch of independent estimates, as shown in Eq. (2). This results in faster convergence than the energy deviations used in sample-based approaches; we have a detailed theoretical and numerical analysis in the SI [SM]. In contrast to the one-by-one generation process of PixelCNN, the NF model is significantly faster due to its all-to-all style of generation. Including training time, the NF model matches Hamiltonian Monte Carlo in estimating $18$ temperatures at a convergence accuracy of $10^{-3}$, and it becomes faster as the number of temperatures increases. Additionally, the low STD allows us to estimate the heat capacity using the first-order derivative, which offers a significant speedup and reduces the memory requirement. We give a detailed description of this in SI [SM]. The proposed method is also effective for the XY model with lattice sizes up to 258 × 258. Larger lattices can be accommodated with increased GPU memory or additional GPU cards.
Conclusion and Outlook
@@
Here, we present a novel method for variational modeling of the canonical ensemble in a continuous temperature range. In the proposed method, any tractable density generative model can be turned into a VaTD model by directly including temperature as a part of the input. By converting the variational definition of Boltzmann distribution into a temperature-dependent loss, we propose the VaTD training framework to unbiasedly optimize the model. This allows thermodynamic quantities, such as free energy and partition function, to be differentiable functions of temperature. Resembling analytic solutions, this differentiable free energy gives a concise way of expressing thermodynamic quantities as its derivatives. Our method is as accurate as the unbiased MCMC method but more efficient. Moreover, as a single model with temperature dependence, its direct-sampling estimation remains valid across a continuous temperature range. In contrast to previous deep generative models that optimize at a single temperature, our method preserves the essential temperature dependence. As a result, there is no need for training additional models at discrete temperature points, and a single model can reliably capture subtle thermal effects, even at PTs.
@@
As a general training framework, with minor modification, we expect to have immediate practical applications by adapting our method to these popular variational deep generative models at a single temperature [equivariantFlowLattice, lattice2, freePart1, freePart2, lwQuantumDots, lwelegas, lw2023PRL, prr]. Moreover, our proposed method can be seen as a generalization of the variational mean-field method, which consists of multiple conditional parameters. While our current implementation focuses on temperature, it opens up intriguing possibilities for exploring other parameters, such as pressure and external magnetic field. In numerical experiments, we mainly applied our method to lattice models, but it is also viable to extend its application to more realistic systems with many meta-stable states, e.g., atomistic simulations.
Acknowledgments
@@ Acknowledgements.
We thank Lei Wang, Wen-Qing Xie, Chu Li, Tao Li, and Jun-Ting Yu for many useful discussions. S.-H.L. and D.P. acknowledges support from Hong Kong Research Grants Council (RGC) (GRF-16302423, GRF-16301723). D.P. acknowledges support from the Croucher Foundation through the Croucher Innovation Award, National Natural Science Foundation of China (NSFC) through the Excellent Young Scientists Fund (22022310), and the NSFC/RGC Joint Research Scheme (N_HKUST664/24). Part of this work was carried out using computational resources from the National Supercomputer Center in Guangzhou, China, and the X-GPU cluster supported by the RGC Collaborative Research Fund C6021-19EF.
Data availability—The code implementation is publicly available at [Code].
Supplementary Information: Deep generative modelling of canonical ensemble with differentiable thermal properties
Appendix A Introduction to tractable density Generative Model
The tractable density generative model is a special kind of generative model. Generative model refers to deep neural network models that can generate samples in a high-dimensional space. This is different from the discriminative model, where usually only a one-dimensional output is given as the category or yes-or-no label of the input data. Examples of the generative model are variational autoencoders (VAE) [vae] and generative adversarial networks (GAN) [gan]. Most of generative models are viewed as probability distributions, and the generation of samples is a one-shot i.i.d. sampling process.
The tractable density generative model is different from general generative models. As an tractable density, it can give the exact probability of the sample. This is usually achieved with specially designed structures or transformations of the neural network model. As two examples, the two most used tractable density generative models are the normalizing flow model [normalizingflow1, normalizingflow2, pwlinearAndquadratic, nsf, cubic] and variational autoregressive models (e.g., PixelCNN [pixelRCNN]). We have introduced the two in the numerical experiments. With this tractable density feature, one can use the model as a sampler inside MCMC algorithm [prl, PREfromRef]. This is not possible with other types of generative models, as the Metropolis-Hasting rule requires knowledge of the proposal probability. Furthermore, tractable density models trained with reverse KLD (Eq. 4) will allow the estimation of partition function and free energy, as proposed in [prl, wuDianPRL].
Appendix B Boltzmann Distribution as the Variational Minimization of the Free Energy
Given a probability distribution $q(\mathbf{x})$ and an energy function $E(\mathbf{x})$, the corresponding free energy is
$F=\langle E\rangle-TS=\sum_{\{\mathbf{x}\}}\frac{1}{\beta}q(\mathbf{x})\log q(\mathbf{x})+q(\mathbf{x})E(\mathbf{x}).$ (S1)
@@ Appendix B Boltzmann Distribution as the Variational Minimization of the Free Energy
Considering the probability distribution $q^{*}$ that minimizes the free energy at all temperatures, then this $q^{*}$ should be the solution to the following optimization problem,
@@
This optimization can be analytically solved via the Lagrange multiplier method [optimbook], which gives the Boltzmann distribution as a function of temperature, i.e.,
@@
Using an tractable density model, one can directly sample $\bm{x}$ to perform statistical averaging as an estimation of the summation over $\{\bm{x}\}$, and the probability normalization condition is automatically satisfied, so we get
$\min_{\theta}\mathop{\mathbb{E}}_{\beta\sim U}\frac{1}{\beta}\Big[\mathop{\mathbb{E}}_{\bm{x}\sim q_{\theta}(\cdot,\beta)}\log q_{\theta}(\mathbf{x})+\beta E(\mathbf{x})\Big],$ (S7)
@@
Appendix C Reparameterization and REINFORCE Estimation
In this section, we introduce the two methods for calculating the derivative of expectation operation. In the main text, these methods are used to compute the optimization gradients and the derivatives of free energy.
C.1 First- and second-order reparameterization
The reparameterization introduces a new random variable $\epsilon$ and rewrites the original expectation into an expectation over the unparameterized distribution of $\epsilon$, i.e.,
@@ C.1 First- and second-order reparameterization
The first- and second-order derivatives can be reformulated as
@@
C.2 First- and second-order REINFORCE estimation
@@
The variance of estimates of Eq. (S11) can thus be reduced via: [reparamAndreinforce]
@@
where $b$ is the baseline. The baseline value $b$ has many different formulas [reparamAndreinforce] and here we used the simplest one:
@@
$+\sum_{\{\mathbf{x}\}}q_{\theta}(\mathbf{x})\nabla_{\theta}^{2}f(\mathbf{x};\theta)+2\nabla_{\theta}f(\mathbf{x};\theta)q_{\theta}(\mathbf{x})\nabla_{\theta}\log q_{\theta}(\mathbf{x})$
@@
The variance reduction method can also be employed as
@@
C.3 Reparameterization of acceptance-rejection sampling
The acceptance-rejection sampling is a common way to draw samples from certain relatively complex distributions, such as the von Mises distribution. The reparameterization of acceptance-rejection sampling is given in Ref. [acptReparam]. For a better illustration, we first give the general framework of acceptance-rejection sampling.
In acceptance-rejection sampling, to sample distribution $q_{\theta}(z)$, we first draw a random variable $\epsilon$ from the distribution $s(\epsilon)$. It is then transformed using a parameterized transformation $h_{\theta}$, similar to the NF, we should consider the Jacobian of this transformation as the change of the probability distribution. Then we note its changed possibility as $r_{\theta}({z})$, where ${z}=h_{\theta}(\epsilon)$. The sample is accepted with the probability $\min\{1,\frac{q_{\theta}({z})}{M_{\theta}r_{\theta}({z})}\}$ or resampled till being accepted, as in Alg. 1.
@@ 5 $u\sim{U}(0,1)$;
$\pi_{\theta}(\epsilon)=$ $\int M_{\theta}s(\epsilon)\mathds{1}\left(0<u<\frac{q_{\theta}(h_{\theta}(\epsilon))}{M_{\theta}r(h_{\theta}(\epsilon);\theta)}\right)du$ (S19)
@@
Adapting the reparameterization, the first-order derivative can be reformulated as
@@
Then, using the log-derivative trick of REINFORCE, this becomes
@@
We generalize the scheme in Ref. [acptReparam], and give the second-order derivative
@@
This can be reformulated as
$\nabla_{\theta}\int s(\epsilon)\frac{q_{\theta}(h_{\theta}(\epsilon))}{r(h_{\theta}(\epsilon);\theta)}\nabla_{\theta}f(h_{\theta}(\epsilon);\theta)d\epsilon$ (S23)
@@
$\mathop{\mathbb{E}}_{\epsilon\sim s(\epsilon)}\nabla_{\theta}\frac{q_{\theta}(h_{\theta}(\epsilon))}{r(h_{\theta}(\epsilon);\theta)}=\int\nabla_{\theta}\pi_{\theta}(\epsilon)d\epsilon=0,$ (S24)
$\mathop{\mathbb{E}}_{\epsilon\sim s(\epsilon)}\nabla_{\theta}^{2}\frac{q_{\theta}(h_{\theta}(\epsilon))}{r(h_{\theta}(\epsilon);\theta)}=\int\nabla_{\theta}^{2}\pi_{\theta}(\epsilon)d\epsilon=0,$
so the variance reduction method [reparamAndreinforce] can also be applied here.
@@
Similar to the NF model, the $r_{\theta}$ distribution can be derived using the $s$ distribution and the Jacobian of $h_{\theta}$.
Appendix D Thermodynamic Quantities from the Differentiation of the Free Energy
D.1 Theoretical analysis of the derivative estimation method compared to the statistical averaging with direct sampling
After training a VaTD model, in addition to the derivative estimation method used in the experiment (as shown in Eq. (2)), an alternative approach to estimating mean energy and heat capacity is to perform statistical averaging with direct sampling, i.e.,
@@ Appendix D Thermodynamic Quantities from the Differentiation of the Free Energy
For a better illustration, we first establish a connection between the two methods.
Using Eq. (2), given the estimated $\log\bar{\mathcal{Z}}(\beta)$, the mean energy is
@@
A similar analysis can be done for the heat capacity with the same conclusion; that is, when the training is perfectly done, the derivative estimation method gives the same results as those obtained from statistical averaging.
@@
For imperfect training, the two methods give different results. The derivative estimation, which is closely related to the free energy (the loss function), typically has better accuracy compared to statistical averaging. Additionally, one can view the VaTD model as a generalization of the variational mean-field method. In the mean-field theory, when the variational ansatz $q_{\theta}$ is not flexible enough, the $q_{\theta}$ does not converge to the exact distribution, but still can have a relatively accurate estimate of the free energy and partition function.
@@
On the other hand, even with perfect or near-perfect training, it is better to choose the derivative estimation method. This is because the derivative estimation method has a much faster convergence speed; from Eq. (4), one can see that a perfect model will give uniformly the same estimate of the free energy for any $\bm{x}$. In this sense, a batch of only one sample will converge.
@@
We further give numerical tests in the following sections to demonstrate these two advantages of the derivative estimation method.
D.2 Theoretical analysis of the convergence time of derivative estimation
In VaTD models, when estimating the mean energy or magnetization, each sample is i.i.d. sampled, so the estimates are also i.i.d., which means the convergence error $\sigma_{N}$ follows the variance of the mean of $N$ independent values each with variance $\sigma^{2}$, i.e.,
$\sigma_{N}=\sqrt{\frac{\sigma^{2}}{N}}$ (S32)
In MCMC, where samples are correlated, the estimates from each sample are also correlated. Then, as a simple analysis, one can use a $\Delta t$ as the time step interval to take samples, which should be much larger than the autocorrelation time $\tau$ to make sure that each sample closes to an i.i.d. sample. This means the effective sample batch size is $N_{eff}=N/\Delta t$, i.e., the autocorrelation time introduces a scaling factor that slows down the convergence. At PT, the correlation effect will dominate, where large autocorrelation times are observed. So, i.i.d. samplers, such as the proposed VaTD model, converge faster than MCMC, especially at PT.
In the estimation of heat capacity, for MCMC, convergence is harder. This is because, the heat capacity is expressed as the deviation of energy in a batch of sampled states, which is a macroscopic statistical quantity and inevitably correlated to all samples in the batch. This means heat capacity will converge slower than, for example, mean energy or magnetization. A simple convergence analysis like Eq. (S32) will not apply. More advanced methods, like the bootstrap method [bootstrap], is required. For tractable density models with only a single fixed temperature, like [prl] and [wuDianPRL], this is also the case, as they’re also sample-based. In contrast, for VaTD models, by making the temperature differentiable, one can estimate the heat capacity via Eq. (2). This is a fundamentally different approach. As in Eq. (2), each i.i.d. sample gives an i.i.d. estimate of the free energy and thus an i.i.d. estimate of the heat capacity. So, the final estimate is still a mean over a batch of i.i.d. estimates, where Eq. (S32) still applies and a faster convergence can be achieved.
In the following section, we provide a convergence time comparison for experimental demonstrations.
D.3 Numerical comparisons of convergence times of mean energy and heat capacity
Temperature ($\beta$)
Methods 0.65 0.70 0.90 0.91 0.95 1.15 1.20 1.25
HMC ($\langle E\rangle$) 497.46 776.07 526.67 518.43 435.43 452.98 469.50 568.46
VaTD ($\langle E\rangle$) 7.03 9.69 38.94 35.98 23.61 8.24 6.70 6.39
HMC ($C_{v}$) 3164.93 3959.81 21863.21 11719.66 19654.64 5575.14 5495.62 4883.55
VaTD ($C_{v}$) 35.98 60.19 256.01 243.72 206.86 146.40 143.21 142.83
Table S1: Comparison of convergence times of Hamiltonian Monte Carlo (HMC) and VaTD model in estimating heat capacity ($C_{v}$) and mean energy ($\langle E\rangle$) of the $16\times 16$ 2D XY model. The convergence threshold is $10^{-3}$ for the mean energy ($\langle E\rangle$) and heat capacity ($C_{v}$) estimations. For the MCMC algorithm, we used a PyTorch implementation of HMC to leverage GPU parallelization and automatic differentiation for force calculations. For the VaTD model, we employed the same model presented in the manuscript. All the tests are conducted using a same NVIDIA GeForce RTX 4090 GPU card. The time is measured in seconds (s).
Figure S1: A comparison of total time consumption between VaTD and MCMC methods on a 2D $16\times 16$ XY model with periodic boundary conditions. The temperature points are equally spaced within the range of $\beta\in[0.5,1.5]$. The convergence accuracy is set at $10^{-3}$. We estimated the RMHMC time consumption curve using the speedup of $3.5$ relative to HMC, which is a upper bound of the reported best-case speedup of $3$.
Figure S2: The mean training time per epoch per sample of 2D XY models as a function of the site number. The $O(N)$ line is estimated using the time data from the training of the $16\times 16$ 2D XY model.
Figure S3: The total time of training and evaluating $10$ temperature points of 2D XY models as a function of the site number. The $O(N)$ line is estimated using the time data from the $16\times 16$ 2D XY model. The convergence error is set to be $10^{-3}$.
In addition to the total time comparison in the main text, we also compared the convergence times of reaching an error of $10^{-3}$ for the VaTD model and Hamiltonian Monte Carlo (HMC). The results are shown in Tab. S1, and the time unit is second (s).
For HMC, we used a batch size of $4000$, and a time step interval $\Delta t$ of $75$. The burn-in step of heat capacity was $1000$ while mean energy used a burn-in step of $20$. The integration algorithm in HMC is the leapfrog with Verlet time step. We also implemented an adaptive integration step size, which will keep the acceptance ratio above $80\%$. For error estimation in HMC, we used the bootstrap method [bootstrap]. For the VaTD model, we used a batch size of $540$. And we used Eq. (2) to estimate mean energy and heat capacity. For error estimation, as analysised above, we used Eq. (S32). This convergence test is done on $8$ different temperatures, as listed in Tab. S1.
These experimental results accurately reflects the theoretical analysis above. Firstly, in the convergence time comparison, VaTD model is generally 109010-90 times faster than HMC. As explained above, this is mostly due to the direct sampling of i.i.d. samples. Then, at the same temperature, a greater speedup from VaTD is observed in the convergence of heat capacity than mean energy. This is because VaTD can gain extra speedups by expressing heat capacity as a derivative of the free energy, as in Eq. (2).
We further investigated the total time consumption of VaTD compared with MCMC methods in Fig. S1. In this figure, the proposed VaTD model included an initial overhead for training time, while HMC started from the origin. The tested temperature points are equally spaced within the range ($\beta\in[0.5,1.5]$), and we use a convergence accuracy of $10^{-3}$. Fig. S1 shows that VaTD scales more efficiently than HMC, implying a cross point is inevitable. In this plot, the cross point happened at about $18$ temperature points. As direct sampling gives i.i.d. samples without the effect of autocorrelation, this cross point will happen earlier at a lower convergence accuracy.
We also estimated the speed of the Riemann manifold HMC (RMHMC) in this comparison. As reported in [RMHMC], the speedup of RMHMC over HMC ranges from $0.7$ to $3$, depending on the specific sampling datasets. We thus estimated the time consumption curve of RMHMC using the speed ratio of $3.5$ relative to HMC, which is an upper bound of the best reported speedup $3$. In this best-case scenario, RMHMC would postpone the crossover by approximately $50$ additional temperature points.
In Fig. S2, we present a scaling test of training time as a function of different lattice sizes. The results show that the training time scales efficiently, following a sub-$O(N)$ curve without significant slowdown. Particularly, we observe a $6$-fold increase in average training time when the input dimension is increased by a factor of $100$. Similarly, in Fig. S3, we also observed a sub-$O(N)$ curve when plotting the total time required for training and evaluating $10$ temperature points. This efficiency is attributed to the use of a CNN-based neural network architecture, which is inherently less sensitive to input dimensions.
D.4 Estimation of the heat capacity using first-order derivative
As in Eq. (2), calculating the heat capacity requires the second-order derivative of the free energy. However, it is known that the second-order derivatives are not memory-efficient with backpropagation automatic differentiation. This can be easily resolved by the forward automatic differentiation. Nevertheless, backpropagation is more widely used, and as in the XY experiment we achieved very low STDs from the free energy estimates, we propose another scheme to estimate the heat capacity using only the first-order derivative.
Consider the first-order derivative of the log-probability,